OpenStudy (anonymous):

Factor completely: 3x^2 - 27

OpenStudy (anonymous):

@MDoodler

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@shrutipande9

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

hey @hartnn

OpenStudy (anonymous):

so i know that i have to factor out the GCF The terms \(\large\cal\color{lime}{3x^2~and~27~have~a~GCF~of~3}\).

OpenStudy (anonymous):

i know that i have to determine also x^2 - 9 is a Difference of Two Squares so 3(x^2-9) but i dont think that three doesnt have a perfect square

OpenStudy (akashdeepdeb):

Hey @Law&Order You can right this as: \[3x^2 - 27\] Now what can you take in common? *This is what I say in my head* { Well, I see a 3 an \(x^2\) and a 27, what can I do here? Idea! 3 factorizes 27, because 3 * 9 = 27 itself. So let me take 3 in common. Then let me see what I can do. } My spidey senses tell me to take 3 in common. So, we get: \[3x^2 - 27 = 3(x^2 - 9)\] *Talking to myself intensifies* { This does not look so right. Maybe I can factorize more, can I, really? OH YES! I CAN! } If you remember, there is an identity for perfect squares: \(a^2 - b^2 = (a+b)(a-b)\) And we can use it over here. So we get, \[3(x^2 - 9) = 3(x-3)(x+3)\] Getting this? :) Do you think we can factorize more?

OpenStudy (anonymous):

oh i think we can factoize more

OpenStudy (anonymous):

but first i have to check my factors

OpenStudy (anonymous):

\(\color{red}{\large\ (3(x + 3)(x - 3) = 3[x^2 - 3x + 3x - 9]\ }\) = 3[x^2 - 3x + 3x - 9] = 3[x^2 - 9] = 3x^2 - 27 \(\color{red}{\huge\ ✓}\)

OpenStudy (anonymous):

so the answer is \(\huge\ 3(x + 3)(x - 3)\)

OpenStudy (akashdeepdeb):

Yes. :)